i've got maths homework from my teacher i want to know:why do you need to have a take away sign in front of the equation whilst subtracting terms e.g -y-4y what does the first take away sign mean?
Found 2 solutions by AnlytcPhil, Eric Johnston:
Answer by AnlytcPhil(1810)

I've got maths homework from my teacher i want to know:
Why do you need to have a "take away" sign in front of the equation whilst
subtracting terms e.g -y-4y what does the first take away sign mean?

It's not an "equation" because it doesn't have an equal sign " = ".  It's
called an "expression". 

That's not a "take away" sign in front.  It's a NEGATIVE sign.

"MINUS ___" means "take away ___" which means "PLUS NEGATIVE ___"

-y-4y means "NEGATIVE y MINUS 4y" which means "NEGATIVE y PLUS NEGATIVE 4y"

That is,

-y - 4y  means

(-y) + (-4y)

The (-y) means (-1y), so you have

(-1y) + (-4y)

We are adding like y-quantities so since (-1) + (-4) = (-5), we have

(-5y)

or 

-5y

"Take away" signs and "negative" signs look just alike but they are different.

If the " - " sign comes in front it's a "negative" sign

If the " - " sign comes between two terms it's a "take away" sign.

Edwin
AnlytcPhil@aol.com

Answer by Eric Johnston(7)

That is a strange question, since when you add the -y to the -4y you obviously get -5y
Are you sure is was not -x-4y ?
It is unusual to put a minus sign in front of the the first variable, so it may be mistake, but it may be intentional. If you have some convenient positive value in the series I would tend put that first to help the user, but sometimes it may be appropriate to put a certain variable first regardless.
For example, you might do it yourself.
Say the problem was 3b -a -5d +2c = Answer
You have a list of values of a, b, c and d and want to write out the equation and work out the answer for each one. It is convenient for you to rearrange the equation thus:
Answer = -a +3b +2c -5d (it tidies up the sequence a,b,c,d so you are less likely to confuse yourself when substituting in the values)
Now when you substitute values for a, you get
a=100 Answer = -100 +3b +2c -5d
a=-10 Answer = 10 +3b +2c -5d
Note carefully how when substituting the -10 it turns into +10
It may be helpful to consider the -a at the front as -1 * a